#26 please, thank you.
Use to solve next two problems: a=r? - sº; b = 2rs ; c=r?+s? 25.) r=1, s=1 26.) r=4, s=3
Answer #1

25) a =p - =a=12 – 12 → a=1-1 a=0 b = 2rs b=2(1)(1) b=2 c=p2 +s? c= 12 +12 c=1+1 c=2

26) a = p? - → a= 42 – 32 > a = 16 - 9 → a=7 b = 2rs b=2(4)(3) b=24 <=p2 +52 =c=42 +32 > c= 16+9 c=25